. A perfectly competitive firm has the cost function TC = 1000 + 2Q + 0.1 Q2. What is the

Lowest price at which this firm can break even?


1
Expert's answer
2021-05-24T13:20:45-0400

"ATC=\frac{TC}{q}"


"TC=1000+2Q+0.1Q^2"

"MC=\frac{\delta TC}{\delta q}=2+0.2Q"


"ATC=\frac{1000+2Q+0.1Q^2}{Q}"


"ATC=\frac{1000}{Q}+2Q+0.1Q^2"


Set MC=ATC

"2+0.2Q=\frac{1000}{Q}+2Q+0.1Q^2"


"0.2Q-0.1Q=\frac{1000}{Q}+2-2"


"0.1Q=\frac{1000}{Q}"


"0.1Q^2=1000\\Q^2=\frac{1000}{0.1}\\Q^2=10,000\\Q=100"


Price=MR=MC

"MC = 2 + 0.2Q"


substituting Q=100,

"MC = 2 + 0.2 (100) = 2 + 20 = 22"


The lowest price at which the firm can break even is 22